# A low complexity contextual stacked ensemble-learning approach for pedestrian intent prediction

### v1

@@ III-C CSE: Contextual Stacked Ensemble
Similar to the approach used in [22] for multimodal feature fusion, we propose a stack-ensemble learning method to fuse features from each of the three models defined above. Each of M1, M2, and M3 is trained by a five-fold stratified cross-validation with a fixed sampling seed 1. The variant of cross-validation ensures each fold preserves the same class ratio. The last fully connected layer of each of the three models M1, M2, and M3 produces an output S1, S2, or S3, respectively (as shown in Figures 5 and 1) when the test input is a five-fold validation dataset. We then concatenated S1, S2, S3 to form the aggregated features that are then fed into a CSE classifier to boost overall prediction. The classifier is then followed by a softmax and an argmax function to generate the final Crossing/Not-Crossing classification results (see CSE model structure in Figure 5).
Figure 5: CSE: Model structure of Contextual Stacked Ensemble learning. In stratified cross-validation, we first train models with five-fold training datasets. An example is shown in Fold 1, where three models are trained on four blocks (see bottom-to-top 4 blocks marked train set Fold 1 above). In this case, the validation block (top purple) is used to test each of the three models. This process is repeated with the iteration of folds; for instance, the next fold would use the green block for validation and train using the remaining blocks. The testing results of all five folds are concatenated to form S1, S2, and S3 as an input of CSE classifier.
@@ IV Experiments
We conducted the experiments of our proposed method on two datasets: (1) Joint Attention in Autonomous Driving (JAAD) dataset[23] as described in Section IV-A. Next, we introduce the data preprocessing, where sequences of frames are processed, and three different features related to pedestrians are extracted. The first is keypoint representation of pedestrians in bbx (Section IV-B1). The second consists of contextual information around this pedestrian including time-series data and categorical data (Section IV-B2). The last is the historical trajectory of the pedestrian, which is extracted by analyzing the changing position of the bounding box with respect to the frame (Section IV-B3).
@@ IV-A Dataset description
In this work, we use the Joint Attention in Autonomous Driving (JAAD) Dataset [23], which is richly annotated and contains $V=346$ video clips (each is around five to ten seconds long) extracted from over $240$ hours of driving footage from dash cameras positioned at the front of the car. The dataset provides frame-wise, distinct labels such as pedestrian reaction toward drivers, pedestrians’ demographic attributes, and traffic conditions, which are not found in other datasets, such as the Multiple Object Tracking (MOT) Challenge  [^2] https://motchallenge.net. There are two versions of the same dataset, JAADbeh and JAADall, where JAADbeh is a subset of JAADall, and the latter includes additional annotations of pedestrians in the background.  [^3] JAADall dataset incorporates all non-labeled pedestrians in JAADbeh by manually annotating these to Not-Crossing class. (see details in Section IV-B).
IV-B Data preprocessing
@@
The first data representation aims to capture the pose of a pedestrian by examining the essential keypoints of the human body following a skeleton-ization approach. The reasoning behind this representation is multifaceted. First, it avoids the capturing of personal information (e.g., face, hair, clothing, etc.) while retaining the essential information for pose analysis and, therefore, action prediction. Second, reducing the complexity of the data representation decreases the computational requirement for classification, which indicates that this task can be conducted at the edge with limited processing power. Third, given that the data is significantly compressed with skeleton-ization, it is then possible to efficiently transfer data to the central server for further processing. We obtain 17 keypoints for each pedestrian by adopting a MMpose  [24] as shown in Figure 6. MMpose is a powerful open source toolbox that provides weights from more than 200 networks for human keypoint detection.
To obtain 17 keypoints, each sequence of $F_{max}$ bbxs represents a data sample (i.e., a pedestrian) and is labeled as Crossing if any bbx found in frame $F_{i}$ ($1\leq i\leq F_{max}$) is categorised as such, otherwise, it is labeled as Not-Crossing.
Figure 6: The MMpose returns $17$ keypoints for each bbx. (Left) Not-Crossing (right) Crossing
IV-B2 Contextual data
In this section, we examine key attributes and investigate if this additional information (discussed in Section I) is more likely to improve the prediction of pedestrians (Crossing/Not Crossing) in the PIP problem. There are three categories of attributes that can be obtained:
  • Pedestrian Behavior shown in Figure 7 and includes Action={Walking, Standing}, Hand Gesture={Greet, Yield, Rightofway, other}, Reaction={Slow-down, Speed-up, Clear-path}, Look={Yes, No}, and Intersection={Yes, No}. All except Intersection are time-series data.
  • Vehicle speed before pedestrian’s crossing action shown in Figure 9
Pedestrians that are not labeled as Crossing or Not-Crossing have been marked as Irrelevant in Figures 78, and 9.
@@ IV-B2 Contextual data
It may be counter-intuitive at first that some of the Pedestrian Behavior attributes shown in Figure 7 do not have a distinct pattern for Crossing or Not-Crossing. For instance, pedestrians that are looking and those that are walking are more likely to cross but the distinction with the Not-Crossing category is minimal. Careful analysis indicates that pedestrians who stand may have the intention to cross and those who look may be standing-by. Reaction shows that Crossing pedestrians are more likely to slow down while Not-Crossing are more likely to clear path. Intersection shows that Crossing pedestrians are likely to stand on the intersection.
Examining the Traffic Annotation attributes can similarly be misleading where the occurrences of the two classes Crossing, and Irrelevant with each of the attributes Road Type or Crosswalk and is almost identical. As such, it is not possible to distinguish between these two classes by examining these attributes.
Pedestrians that are labeled as Irrelevant are more likely to be seen in the traffic scene when the traffic light is green than the other two classes. Following careful investigation of such occurrences, we noted in most cases the pedestrians are not on the same lane as the ego-vehicle. Instead, they could be on the zebra crossing parallel to the ego-vehicle’s lane or they could be on the sidewalks.
@@ IV-B3 Pedestrian trajectory
The trajectory of the pedestrian within each given frame carries contextual information that can be useful in solving the PIP problem. The location of the pedestrian is equated to the relative position of bbx with respect to the frame. In this work, we calculate the relative position of the central bbx pixel located at $[W_{c}=W/2,H_{c}=H/2]$, where $W$ and $H$ are the width and height of the bbx. The pedestrian’s depth, $Z$, is inferred from the size of the bbx where a small value indicates that the bbx is far away from the camera and vice-versa. Knowing that the absolute coordinates $(W_{c},H_{c})$ and the size $Z$ can be biased by different frame sizes, we normalize the values such that the pedestrian normalized location is $(W_{c}/W_{F},H_{c}/H_{F})$ and the normalized size is $Z/(H_{F}\times W_{F})$, where $H_{F}$ is the height of the frame and $W_{F}$ is the width. In addition, we calculate the change of normalized size over time $\Delta Z/(H_{F}\times W_{F})$ as well as the speed of movement, which was estimated by getting the ratio of the normalized traveled distance over time.
For each video clip $\mathbb{S}_{v}$, the normalized positioning vector $q_{v,i}=[{W_{c}}_{i}/W_{F},{H_{c}}_{i}/H_{F},Z_{i}/(H_{F}\times W_{F}),\Delta Z/(H_{F}\times W_{F}),D/t]$ of each bbx containing a pedestrian in frames $F_{i}$ ($i=[0,\dots,F_{max}$]) is calculated. Thus, the trajectory of a pedestrian in video clip $\mathbb{S}_{v}$ is represented by $\boldsymbol{\tau_{v}}$.
@@ IV-C Results and Discussion
We address the problem of pedestrian intention prediction, predicting label $\hat{y}$ for each given input sliding window $\mathbb{W}_{v}$ (32 frames) sampled from each video clip $\mathbb{S}_{v}$. In preparation for the model training and testing, we split the dataset to train/val/test subsets based on JAAD’s suggestion. The JAAD’s train and validation were then merged before splitting again for stratified 5-fold cross-validation [26]. For each fold, the method preserves the percentage of Crossing and Not-Crossing data points instead of random sampling on the positive ($y=1$, Crossing) and negative ($y=0$, Not-Crossing, Irrelevant) samples pool. This approach ensures that the training performance of each fold is not biased by a random ratio of classes and preserves all information in the samples. The stratified 5-fold crossing validation was used only in training models on JAADbeh. This method preserves the class ratio and therefore, ensures the training stability, which works well for JAADbeh. However, with an extreme class distribution, such as JAADall, we find that models perform better when classes are balanced. Therefore, we create a customized 5-fold dataset in which each fold has all positive samples from the merged data and randomly selected negative samples (without replacement) from the merged data to match the number of positive samples. Nonetheless, the test results are derived using the test dataset defined in  [^4] https://github.com/ykotseruba/JAAD/tree/JAAD_2.0/split_ids/default for both JAADbeh and JAADall.
TABLE II: Performance of models in TTC (Time to Cross)
Dataset JAADbeh JAADall
Avg TTC 0-2sec 2-5sec 5+sec Overall 0-2sec 2-5sec 5+sec Overall
M1 74 81 85 80 74 74 13 58
M2 89 94 100 94 60 53 68 59
M3 70 82 100 85 79 70 77 75
M1+M2 77 84 93 84 75 72 13 57
M1+M3 82 86 93 87 76 74 13 58
M2+M3 85 92 100 92 75 64 75 70
CSE (M1+M2+M3) 78 84 92 84 69 65 13 52
@@
Each of the three data streams described in Section IV-B is used independently to train three different models, M1, M2, and M3, as shown in Figure 1. An ensemble learning approach is built based on the feature representation of these models to form the proposed CSE. We first present the results of each model in isolation in addition to the CSE in Table I and compare the performance and complexity of these to the state-of-the-art models. As seen in Table I, our proposed CSE using M1, M2, and M3 delivers the best F1 score, outperforming the next best models, Pedgrapgh+ [7] and Faster PCPNet [10], by $3\%$ at a much lower computational complexity. Our model requires $331$ times less FLOPS than Pedgraph+ and $10000$ less than Global PCPA; moreover, our model consists of $5.2$ times less trainable parameters than Pedgraph+ and $4500$ times less than Global PCPA. Recall is an important metric as it reflects the ability of the model to detect Crossing; our model ranks second to Global PCPA [25] with $4\%$ reduction in recall and $99.99\%$ reduction in FLOPS and $99.97\%$ reduction in the number of trainable parameters.
Moreover, we examine the Time to Cross (TTC) performance of the proposed individual models M1, M2, and M3, and their respective combinations using the CSE approach. The results are shown in Table II. It should be clarified that these results reflect the performance of each of these models in predicting the Crossing behavior only in contrast with the results in Table I which represent both classes. It is thus expected that the TTC results of JAADbeh (a subset of JAADall containing all pedestrians with Crossing behavior) are far better and more stable than JAADall. Indeed, increasing the number of frames from $60$ (i.e. $2$ seconds as the video is recorded at $30$ frames per second) to more than $150$ frames ($5+$ seconds) consistently improves the accuracy of predicting Crossing for each of the models M1, M2, M3, and their respective combinations, reaching perfect prediction with $100\%$ accuracy. On the other hand, the results with JAADall which constitutes a majority of non-behavior, i.e., Not-Crossing, exhibit a deteriorating ability to predict Crossing as the number of frames is increased.
The difference in performance between both datasets models is further confirmed by examining the confusion matrices of the CSE approach with M1, M2, and M3 listed in Table I. The results for JAADbeh and JAADall are shown in Figures 10 and 11, respectively. The proposed model for JAADbeh is very good at predicting Crossing ($80\%$ accuracy) but fails to learn the Not-Crossing features (see Figure 10). It is thus, intuitive that the TTC improves with an increased number of frames for the JAADbeh, as seen in Table II). On the other hand, the performance of the JAADall model exhibits the opposite trend with $89\%$ accuracy in predicting Not-Crossing and poor performance for Crossing predictions (see Figure 11).
Figure 10: CSE confusion matrix for JAADbeh
Figure 11: CSE confusion matrix for JAADall

### v2

@@ III-C CSE: Contextual Stacked Ensemble
Similar to the approach used in [21] for multimodal feature fusion, we propose a stack-ensemble learning method to fuse features from two or three models defined above. Each of M1, M2, and M3 is trained by a five-fold stratified cross-validation with a fixed sampling seed 1. The variant of cross-validation ensures each fold preserves the same class ratio. The last fully connected layer of each of the three models M1, M2, and M3 produces an output S1, S2, or S3, respectively (as shown in Figures 5 and 1) when the test input is a five-fold validation dataset. We then concatenated S1, S2, S3 (or a subset) to form the aggregated features that are then fed into a CSE classifier to boost overall prediction. The classifier is then followed by the sigmoid and argmax functions to generate the final Crossing/Not-Crossing classification results (see CSE model structure in Figure 5).
Figure 5: CSE: Model structure of Contextual Stacked Ensemble learning. In stratified cross-validation, we first train models with five-fold training datasets. An example is shown in Fold 1, where three models are trained on four blocks (see bottom-to-top 4 blocks marked train set Fold 1 above). In this case, the validation block (top purple) is used to test each of the three models. This process is repeated with the iteration of folds; for instance, the next fold would use the green block for validation and train using the remaining blocks. The testing results of all five folds are concatenated to form S1, S2, and S3 (or a subet) as an input of CSE classifier.
@@ IV Experiments
We conducted the experiments of our proposed method on two different labeled versions of the Joint Attention in Autonomous Driving (JAAD) dataset: JAADbeh[22] and JAADall, which includes additional annotations of pedestrians as described in Section IV-A. The rest of this section discusses data engineering, where a sequence of $F_{max}$ frames (defined by the sliding window) is processed, and three different groups of features are extracted. Firstly, the keypoint representation of labeled pedestrians bounded by bbx is discussed in Section IV-B1. Secondly, extracting contextual information around pedestrians, including time-series and categorical data, is described in Section IV-B2). The third group of features describes the historical trajectory of the pedestrian, which is extracted by analyzing the changing position of the bbx with respect to the frame (Section IV-B3).
@@ IV-A Dataset description
In this work, we use the JAAD Dataset, which is richly annotated and contains $V=346$ video clips (each is around five to ten seconds long) extracted from over $240$ hours of driving footage from dash cameras positioned at the front of the car. The dataset provides frame-wise, distinct labels including pedestrian reaction toward drivers, pedestrians’ demographic attributes, and traffic conditions, which are not found in other datasets, such as the Multiple Object Tracking (MOT) Challenge  [^2] https://motchallenge.net. There are two versions of the same dataset, JAADbeh and JAADall, where JAADbeh is a subset of JAADall, and the latter includes additional annotations of pedestrians in the background.  [^3] JAADall dataset incorporates all non-labeled pedestrians in JAADbeh by manually annotating these to Not-Crossing class. (see details in Section IV-B).
IV-B Feature Engineering
@@ IV-B1 Skeleton-ization
The first data representation aims to capture the pose of a pedestrian by examining the essential keypoints of the human body following a skeleton-ization approach. The reasoning behind this representation is multifaceted. First, it avoids the capturing of personal information (e.g., face, hair, clothing, etc.) while retaining the essential information for pose analysis and, therefore, action prediction. In this work, we first apply the MMpose [^4] https://github.com/open-mmlab/mmpose/tree/main algorithm which returns $17$ key points. Next, we combine the keypoints representing the left eye, right eye, and nose into a single keypoint that represents the head. Similalry, we replace the keypoints representaing the left and right shoulders with a single point in the middle that represents the thorax. It follows that our M1 model is based on $14$ keypoints instead of $17$.
Figure 6: The MMpose returns $17$ keypoints for each bbx. (Left) Not-Crossing (right) Crossing . Note: In this work, the $17$ keypoints are further reduced to $14$ (see Section IV-B1).
@@
IV-B2 Contextual data
In this section, we explore how well contextual features particular to JAAD can represent pedestrians’ intentions. There are three categories of attributes that can be obtained:
  • Pedestrian Behavior shown in Figure 7. It includes Action={Walking, Standing}, Hand Gesture={Greet, Yield, Rightofway, other}, Reaction={Slow-down, Speed-up, Clear-path}, Look={Yes, No}, and Intersection={Yes, No}. All except Intersection are time-series data.
  • Vehicle speed before pedestrian’s crossing action shown in Figure 9
Pedestrians that are not labeled as Crossing or Not-Crossing have been marked as Irrelevant in Figures 78, and 9. Overall, some attributes play a vital role in predicting labels that signify a pedestrian’s intention to either Crossing or Not-Crossing. In the following section, we explore how these features contribute to label prediction.
It may be counter-intuitive at first that some of the Pedestrian Behavior attributes shown in Figure 7 do not have a distinct pattern for Crossing or Not-Crossing. This is particularly evident in the case of pedestrians’ intentions, where those who are looking and those who are walking show a near-equal likelihood to Crossing or Not-Crossing. However, careful analysis indicates that pedestrians who standare more likely to intend to cross, and those who look may be standing-by. Reaction shows that Crossing pedestrians are more likely to slow down while Not-Crossing are more likely to clear path. Intersection shows that Crossing pedestrians are likely to stand on the intersection.
While examining the Traffic Annotation attributes, it seems road type and crosswalk are similar for both the Crossing and Irrelevant labels, which indicates that they carry weak representation of pedestrian intent. Moreover, pedestrians labeled as Irrelevant are more likely to be seen in the traffic scene when the traffic light is green than the other two classes. Following a careful investigation of such occurrences, we noted in most cases, pedestrians are not in the same lane as the ego-vehicle. Instead, they could be on the zebra crossing parallel to the ego-vehicle’s lane or on the sidewalks.
@@ IV-B3 Pedestrian trajectory
The trajectory of the pedestrian within each given sliding window carries contextual information that can be useful in solving the PIP problem. The location of the pedestrian is equated to the relative position of bbx with respect to the frame. In this work, we calculate the relative position of the central bbx pixel located at $W_{c}=W/2,H_{c}=H/2]$, where $W$ and $H$ are the width and height of the bbx. The pedestrian’s depth, $Z$, is inferred from the size of the bbx where a small value indicates that the bbx is far away from the camera and vice-versa. Knowing that the absolute coordinates $(W_{c},H_{c})$ and the size $Z$ can be biased by different frame sizes, we normalize the values such that the pedestrian normalized location is $(W_{c}/W_{F},H_{c}/H_{F})$ and the normalized size is $Z/(H_{F}\times W_{F})$, where $H_{F}$ is the height of the frame and $W_{F}$ is the width. In addition, we calculate the change of normalized size over time $\Delta Z/(H_{F}\times W_{F})$ as well as the speed of movement, which was estimated by getting the ratio of the normalized traveled distance over time.
For each sliding window with index $s\in\{1,\dots,M_{v}/F_{max}\}$, the normalized positioning vector $q_{s,i}=[{W_{c}}_{i}/W_{F},{H_{c}}_{i}/H_{F},Z_{i}/(H_{F}\times W_{F}),\Delta Z/(H_{F}\times W_{F}),D/t]$ of each bbx containing a pedestrian in bbx $F_{i}$ ($i=[0,\dots,F_{max}$]) is calculated. Thus, the trajectory of a pedestrian in sliding window with index with index $s$ is represented by $\boldsymbol{\tau_{s}}$.
@@ IV-C Results and Discussion
We address the problem of PIP for each given input sliding window sampled from each video clip $\mathbb{S}_{v}$. The predicted label $\hat{y}$ is a binary variable, where $\hat{y}=1$ for Crossing if a pedestrian is predicted to intend to cross a street or $\hat{y}=0$ for (Not-Crossing). In preparation for the model training and testing, we split the dataset into a training, validation, and test subsets based on JAAD’s suggestion (see  [^5] https://github.com/ykotseruba/JAAD/tree/JAAD_2.0/split_ids/default). The JAAD’s train and validation were then merged before splitting again for stratified 5-fold cross-validation [24]. For each fold, the method preserves the percentage of Crossing and Not-Crossing data points instead of random sampling on the positive ($y=1$, Crossing) and negative ($y=0$, Not-Crossing, Irrelevant) samples pool. This approach ensures that the training performance of each fold is not biased by a random ratio of classes and preserves all information in the samples. The stratified 5-fold crossing validation was used only in training models on JAADbeh. This method preserves the class ratio and therefore, ensures the training stability, which works well for JAADbeh. However, with an extreme class distribution, such as JAADall, we find that models perform better when classes are balanced. Therefore, we create a customized 5-fold dataset in which each fold has all positive samples from the merged data and randomly selected negative samples (without replacement) from the merged data to match the number of positive samples. Nonetheless, the test results are derived using the test dataset defined in Footnote 5 for both JAADbeh and JAADall in order to enable fair comparison with prior art. Moreover, it is worth noting that there exists an inconsistent class distribution between the train and test set, especially for JAADall. In JAADall, the imbalance ratio of Crossing::Not-Crossing is $1:1$ in the train set but $7:93$ in the test set. On the other hand, in JAADbeh, it is $56:44$ in train set but $4:6$ in the test set.
Each of the three data streams described in Section IV-B is used independently to train three different models, M1, M2, and M3, as shown in Figure 1. The CSE approach shown in 5 is applied to each combination of two models (2-model-based CSE) and to the three models (3-model-based CSE). We first present the results of each independent model in addition to the different CSE options in Table I and compare the performance and complexity of these to the state-of-the-art models.
@@
Our results are compared to the state of the art in both performance and complexity including [23], [6] and [9] which were tested using JAADall. We also compared our models with [9] using JAADbeh. As seen in Table I, our proposed CSE using M1, M2, and M3 delivers the best F1 score, outperforming the next best models, Pedgrapgh+ [6] and Faster PCPNet [9], by $3\%$ at a much lower computational complexity. Our model requires $331$ times less FLOPS than Pedgraph+ and $10,000$ less than Global PCPA; moreover, our model consists of $5.2$ times less trainable parameters than Pedgraph+ and $4500$ times less than Global PCPA. Recall is an important metric as it reflects the ability of the model to detect Crossing; our model ranks second to Global PCPA [23] with $4\%$ reduction in recall and $99.99\%$ reduction in FLOPS and $99.97\%$ reduction in the number of trainable parameters.
TABLE II: JAADbeh: Comparative analysis of (A) baseline model (M2) and CSE(M2+M3), (B) baseline model CSE(M2+M3) and additional model M1. The (baseline) models are shown between parenthesis.
JAADbeh
(A) (B)
Baseline\(Baseline)+additional (M2)+M3 Baseline\(Baseline)+additional
(M2+M3)
+M1
(M2) false predictions 0.4 (M2+M3) false predictions 0.35
M3 false predictions 0.39 M1 false predictions 0.42
Common predictions 0.63 Common predictions 0.66
Common false predictions 0.21 Common false predictions 0.21
Actual false predictions 0.35 Actual false predictions 0.40
Union of false predictions 0.58 Union of false predictions 0.56
Correlation coefficient
of M2 and M3
confidence
0.41
Correlation coefficient
of (M2+M3) and M1
confidence
0.28
TABLE III: JAADbeh: Comparative analysis of (A) baseline model (M1) and CSE(M1+M3), (B) baseline model CSE(M1+M3) and additional model M2. The (baseline) models are shown between parenthesis.
JAADall
(A) (B)
Baseline\(Baseline)+additional (M1)+M3 Baseline\(Baseline)+additional
(M1+M3)
+M2
(M1) false predictions 0.19 (M1+M3) 0.17
M3 false predictions 0.32 M2 0.36
Common predictions 0.68 Common predictions 0.59
Common false predictions 0.1 Common false predictions 0.06
Actual false predictions 0.17 Actual false predictions 0.12
Union false predictions 0.41 Union false predictions 0.47
Correlation coefficient
of M1 and M3
confidence
0.284
Correlation coefficient
of (M1+M3) and M2
confidence
0.06
IV-C1 Sensitivity analysis
in this section, we investigate the causes for the differing CSE (M1+M2+M3) gain seen in Table I between JAADbeh and JAADall. To this end, we study the false predictions of (A) the best performing single model in comparison with the best 2-model-based CSE approach (Table II Part (A) for JAADbeh and Table III Part (A) for JAADall) and (B) the best 2-model-based CSE approach in comparison with the 3-model-based CSE approach (Table II Part (B) for JAADbeh and Table III Part (B) JAADall). As shown in Table I, M2 has the best single-model performance (the baseline model) on JAADbeh whereas CSE(M2+M3) is the best 2-model-based CSE approach. As such, in Table II, M1 is the baseline model compared to CSE(M2+M3) (M3 is the additional model) in Part (A), whereas in Part (B), CSE(M2+M3) becomes the baseline compared with CSE(M1+M2+M3) (M1 is the additional model in this case). Similarly, M1 is the best single performing model thus, is considered the baseline for Part (A) of Table III with M3 as additional model; moreover, CSE(M1+M3) is baseline in Part (B) with M2 is additional.
Ensemble learning has proved to be effective if training data and submodels are diverse, meaning model outputs are preferably independent or negatively correlated [25]. Therefore, to examine the input (data) diversity of a merged CSE model, we first examine the output of the baseline and additional models and extract the distribution characteristics listed in Tables II and  III. Firstly, we calculate the ratio of “Common Predictions” between each baseline and additional model which is defined as the number of samples in the test dataset assigned the same class by both. For instance, $63\%$ of the predicted classes ($\hat{y}$) using M1 (baseline) and M3 (additional) in Table II Part (A) are common to both models. Similarly, the ratio of “Common False Predictions” indicates the percentage of the predicted samples that are commonly misclassified by both models; in Table II Part (A), this is $21\%$. Next, we calculate the ratio of “Union of False Predictions” from the baseline and additional models which represents the cardinality of the union of misclassified samples from both models divided by the size of the test dataset; in Table II Part (A), this is $58\%$ of the test dataset. We then calculate the “Actual False Predictions” for CSE models (Table II Part (A), this is $35\%$) and the Pearson product-moment correlation coefficients between the baseline and additional models (Table II Part (A), this is $41\%$). (Note: all scores are normalized by the number of samples in the test set, except for the correlation coefficient). The confusion matrices of the CSE models in Parts (A) and (B) of Tables II and  III are shown in Figures 10 and 11, respectively, to facilitate the analysis.
In Table II Part (A), M2 and M3 have a high ratio of common false predictions, union of false predictions, and high correlation coefficients in their confidence outputs. In addition, CSE(M2+M3) and M1 also have high ratios in these three scores. In this case, the baseline model might not again much in performance when merged with an additional model using the CSE approach. On the other hand, in Table III, despite having non-negligible ratio of union of false predictions, M1 and M3 have a low common false predictions ratio and low correlation coefficient in their confidence outputs. The same applies to CSE(M1+M3) and M2 in Part(B). This means that the baseline and additional models benefit from their combination, making ensemble learning more effective, compared to TableII.
Overall, in Table II, we observe a high correlation coefficient in confidence between the baseline and the additional models, compared to III; in this case, a baseline model might not gain in performance from merging with an additional model. Moreover, we suspect that if both models have a high ratio of common false predictions, merged performance is likely to get worse compared to pre-merged performance. This can be seen by comparing the performance of M2 and CSE(M2+M3), as well as CSE(M2+M3) and M1 in Table II. In Table III, however, JAADall has a lower correlation coefficient and common false predictions ratio even after the first merging. We conclude that it is the reason why the CSE(M1+M2+M3) results for JAADall reported in Table I and confirmed by the confusion matrices in Figures 11 are generally better than the results obtained with JAADbeh (see Figures 10).
(a) CSE(M2+M3)
(b) CSE(M1+M2+M3)
Figure 10: Confusion matrices for CSE(M2+M3) and CSE(M1+M2+M3) in JAADbeh.
(a) CSE(M1+M3)
(b) CSE(M1+M2+M3)
Figure 11: Confusion matrices for CSE(M1+M3) and CSE(M1+M2+M3) in JAADall
IV-C2 Generalization analysis
the proposed CSE model is built on a stacked ensemble approach that leverages multimodalities extracted from video footage. M2 learns the intent of pedestrians based on features that are particular to the JAAD dataset, i.e., Pedestrian Behavior, Traffic Background, and Vehicle Speed (see Figures 78, and 9). Due to the lack of such features in other PIP datasets, these cannot further validate the CSE(M1+M2+M3). On the other hand, M1 and M3 learn the intent of pedestrians based on the dynamics of pedestrian key-points during the time window and the relative location of the bounding box, respectively. It follows that both M1 and M3 can be readily applied to other PIP datasets to verify the generalize-ability of the proposed models. Indeed, a stacked ensemble model based on M1 and M3 was used for the IEEE Intelligent Transportation Systems Society student competition in Pedestrian Behavior in 2023, centered on the PSI dataset [26] and secured the first prize. The obtained results using PSI and JAAD for training and the PSI test dataset determined by the competition are summarized in Table IV. As can be seen, the stacked ensemble learning model improves the performance of individual models M1 and M3, as for the JAADall dataset. These results demonstrate the generalization of the CSE approach, albeit without the M2 model.
TABLE IV: Results of M1, M3, and CSE(M1+M3) models with PSI dataset (JAADbeh and JAADall results from Table I repeated for ease of comparison)
JAADbeh JAADall PSI
ACC AUC F1-score ACC AUC F1-score ACC AUC F1-score
M1 58% 65% 58% 81% 85% 61% 64% 71% 61%
M3 61% 69% 61% 68% 79% 52% 60% 65% 57%
CSE(M1+M3) 60% 70% 59% 83% 87% 63% 68% 72% 63%